# A Reanalysis of the FDA’s Benefit-Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine: For 18–25-Year-Old Males, Risks Exceeded Benefits Relative to Hospitalizations

### v1

@@ 2.2. The FDA’s mRNA-1273 Benefits Assessment
$\bullet$ The FDA assumed in its most likely scenario (as well as Scenarios 5, and 6) that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months. One way to gauge the degree of implausibility of the FDA’s case-rate projection is to consider the corresponding projected number of infections. Throughout the pandemic, the number of COVID infections has always been significantly larger than the corresponding number of reported cases. For the period February 2020 through September 2021, the CDC estimated the number of infections to be four times the corresponding number of reported cases [26]. Moreover, the case-to-infection multiplier likely increased (e.g. owing to increased home testing) during the Omicron-dominant period in the U.S., which started mid-December 2021. If we assume that CDC’s 4:1 infection:case ratio holds over the evaluation period, we find the FDA’s evaluation-period case-count projection of 274,947 per 1 million unvaccinated 18–25-year-old males (Scenarios 1, 5, 6) corresponds to 4 · 274,947 = 1,099,788 infections, well over 1 million. Thus, the FDA’s case-total assumption is essentially equivalent to the assumption that among one million unvaccinated males, there would be, over 5 months, 1,000,000 infections. Having well over one million infections in a group of 1 million 18–25 year-olds over 5 months is implausible. Multiple infections over a five-month period would be unlikely [^4] The study [37] conducted over the period 11/1/21 – 4/30/22 provides evidence that multiple evaluation-period Omicron infections were rare: in the study there were 11,307 positive Omicron tests and none was a second positive Omicron test separated by 90 days from another such positive test. given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [38]). [^5] The preprint [38], posted 6/10/21, was later published [39]. Even direct inoculation with SARS-CoV-2 does not necessarily produce infection: in the challenge study [40] of the 36 inoculated, 16 “remained uninfected.” Thus, for 3 of the FDA’s 5 Omicron-based scenarios, including its most likely scenario, the FDA essentially assumed that every unvaccinated male 18–25 would be infected by COVID over the 5-month evaluation period.
@@ 2.3. The FDA’s mRNA-1273 Risks Assessment
On the risk side, the FDA relied on its internal databases for the VAM/P case rates appearing in Table 1 above while using the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of VAM/P cases requiring hospitalization [9, Section 2.3.2].
The FDA underestimated VAM/P risks for mRNA-1273:
@@
$\bullet$ The FDA omitted dose-one VAM/P events. For its mRNA-1273 analysis [9], the FDA computed the evaluation-period benefits of two doses of mRNA-1273 without accounting for the VAM/P risk associated with dose 1, potentially biasing the FDA’s analysis in favor of vaccination. [^6] The FDA explains its decision to exclude VAM/P events for dose 1 of mRNA-1273 as follows: “The risk post dose 1 was ignored since CDC has suggested people who have myocarditis after dose 1 take precaution to receive dose 2, and the majority of myocarditis cases were reported post dose 2” [9, Section 2.3.1]. In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [8, Section 2.3.1]. Hence, for the BNT162b2 analysis, the FDA did account for VAM/P risk associated with dose 1 but did not assess the benefits of dose 1 alone during the interval starting with dose 1 and ending at the point of full vaccination (7 days after dose 2), potentially biasing the analysis against vaccination. Given that adverse events of vaccination other than VAM/P possibly resulting in hospitalization were ignored by the FDA and that mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit-risk analysis assessing benefits only postdose 2, the choice to include VAM/P risk associated with dose 1 is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in its benefit-risk analysis for mRNA-1273 described in its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 2.4. Thus, we include VAM/P hospitalization risk for dose 1 of mRNA-1273 in our main analysis. In Supplement 2, we conduct a sensitivity analysis showing that accounting for dose 1 benefits during the period leading up to the point of full vaccination does not substantially change our benefit-risk conclusions.
$\bullet$ The FDA ignored evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2. In a “Review Memorandum Addendum” [41], dated 11/19/21, relating to the emergency-use authorization of a booster dose of mRNA-1273, the FDA describes multiple sources of evidence indicating that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2:
@@
[E]vidence was communicated from other regulatory agencies, including agencies from Canada, France, Germany, Norway, and Japan that suggested that there was a higher rate of myocarditis/pericarditis observed in males under 40 years of age following the second dose of the Moderna COVID-19 Vaccine than with the Pfizer-BioNTech COVID-19 Vaccine. For example, Ontario, Canada reported a higher rate of myocarditis in individuals following the Moderna COVID-19 Vaccine than with the Pfizer-BioNTech COVID-19 Vaccine, particularly in males ages 18–24 years ([link provided to Ontario data, later analyzed by Buchan et al. [42]]). …[T]he most recent data presented by the Center for Disease Control and Prevention (CDC) analyzed from the Vaccine Safety Datalink (VSD) indicates a higher rate of myocarditis with the Moderna COVID-19 Vaccine than the Pfizer-BioNTech COVID-19 Vaccine ….
@@
The “most recent data” from the VSD to which the FDA refers was presented at the CDC’s October 2021 Advisory Committee on Immunization Practices (ACIP) meeting [10]. Adjusted-rate ratios for 18–39 year-olds reported in the table on page 25 of [10] suggest that the 2nd-dose VAM/P rate for Moderna’s mRNA-1273 is 2.72 times that of Pfizer’s BNT162b2 and that the corresponding ratio for the 1st dose is 2.63. For 18–39-year-old males, the tabulated data on page 27 of [10] suggests a 2nd-dose rate ratio of 2.26, and, if we assume the same percentage decrease for the 1st dose, then that rate ratio would be 2.19 instead of 2.63. Data from other sources suggests higher rate ratios, especially for dose 2. For example, Buchan et al. found a 2nd dose rate ratio of 5.1 for males aged 18–24 [42, p. 5]. [^7] The specific data (299.5 per million second doses of Moderna vs. 59.2 per million of Pfizer) is the same in the publication [3] and its preprint [42] (available 12/5/2021). A study by Patone et al. [43] of myocarditis cases requiring hospitalization within a 28-day risk interval following a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and approximately 12 excess myocarditis hospitalizations per 1 million 1st doses, with corresponding rates for BNT162b2 being 12 per million 2nd doses and 3 per million 1st doses [43, p. 4].888 Patone et al. [43] had a study population of over 42 million and a study period from 12/1/20 – 11/15/21. The published version [6] of the preprint [43] incorporates an additional month of data and suggests for males $<40$ a rate of 97 myocarditis hospitalizations per million 2nd doses of mRNA-1273 and a rate of 14 per million 1st doses with corresponding BNT162b2 rates being 11 and 4 [6, Table 4].
@@
Despite all the evidence reviewed above, the FDA chose for its mRNA-1273 risk assessment [9] to rely on “preliminary findings” based on its internal databases suggesting “No clear difference in the rate of myocarditis/pericarditis [mRNA-1273 vs. BNT162b2]” [41, p. 2]. The “preliminary findings” to which the FDA refers have “large uncertainty” owing to “small numbers of observed events” [44, p. 16]. [^9] As noted in [44, p. 7] and in [9, Section 2.3.1], VAM/P-data-collection cutoff points, for the four databases on which the FDA relied for its mRNA-1273 risk assessment were: 8/21/21, 7/10/21, 7/31/21, and 6/30/21, each date being over four months before the time of the assessment. Based on these preliminary findings as well as ignoring dose-1 events, the FDA suggests in its mRNA-1273 assessment that for 18–25-year-old males VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [9, Table 4] vs. 131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [8, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations. Throughout this paper, we assume that VAM/P rates for males 18–24 well approximate those for males 18–25.
Based on data from the preceding two paragraphs, we compute a VAM/P-rate estimate for mRNA-1273 using the FDA’s corresponding estimate for BNT162b2 in [8] and VSD VAM/P rate ratios [10]. FDA data available in October 2021 reports 5 myo/pericarditis events postdose 1 of BNT162b2 and 36 postdose 2 [44, p. 14]. Assuming the same ratio of dose 1 and dose 2 events is reflected in the FDA’s projected 131 VAM/P cases for Pfizer’s BNT162b2 [8, Table 3], we find the corresponding estimate for VAM/P cases resulting from fully mRNA-1273 vaccinating 1 million males 18–25 to be
[…]
where we have applied the VSD-based comparison of VAM/P rates, mRNA-1273 vs. BNT162b2, discussed above. The first summand on the left of Equation (2.3.1) represents a VAM/P incidence of 35 cases per million 1st doses of mRNA-1273 among males 18–25, while the second represents approximately 260 cases per million 2nd doses. Using the FDA’s VAM/P case-hospitalization rate of 86% [9], we find the overall case rate of 295 per million full vaccinations corresponds to a rate of approximately 254 VAM/P hospitalizations per million full vaccinations and 224 per million 2nd doses. Thus, the FDA’s estimate of 110 VAM/P hospitalizations per million 2nd doses of mRNA-1273 (used in Scenarios 1–4) is less than half the estimate of 224 per million 2nd doses derived from the FDA’s BNT162b2 estimate and VSD data.
$\bullet$ Internal FDA databases yielded VAM/P rates for mRNA-1273 substantially lower than those found in important studies.    The FDA’s internal databases not only yielded VAM/P-rate estimates for mRNA-1273 inconsistent with the agency’s BNT162b2 estimates and VSD data but also the estimates are substantially lower than those of major international studies by Buchan et al. [42] and Patone et al.  [43]. Furthermore, a U.S. study by Sharff et al. [45] found, with a wide confidence interval, a myo/pericarditis incidence rate of 537.1 per million second doses of an mRNA-vaccine for males ages 18–24 years. The conclusion of [45] is noteworthy: [^10] The preprint [45] and the later publication [4] share the same conclusion.
We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm, which depends on select hospital discharge diagnosis codes. The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees. The VSD should validate its search algorithm to improve its sensitivity for myopericarditis.
We now compare the FDA’s mRNA-1273 projected VAM/P case and hospitalization rates to those of the other sources discussed above.
{adjustwidth}
-1.55cm
@@
Table 3. Estimated VAM/P Events Per Million 2nd Doses of mRNA-1273 Among Males 18–24 or 18–25 years old.* Based on Data Available Before 1 January 2022
@@
* For discussions of the data in columns 3–5, see Section 6 of Supplement 1; in particular, data in column 4 are estimates based on findings from [43] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [42]. Data in column 6 is discussed above in the paragraph of Equation (2.3.1), with 22 being the number of Moderna myo/pericarditis cases contributing to the head-to-head comparison of [10, p. 25].
** All cases of myo/pericarditis contributing to Buchan et al.’s study were judged to meet the Brighton Collaboration’s case definitions (levels 1–3). Applying the Collaboration’s requirements reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [42, p. 5].
*** For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [8, Section 2.3.2.2]. For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [9, Section 2.3.2.2]. We apply the 86% rate in columns 4–6 (which, for column 5, matches the case hospitalization rate for the 7 males 18–24 having myo/pericarditis after dose 2 of an mRNA COVID vaccine: 6 of the 7 were hospitalized). For column 3, Buchan et al. provide an overall case-hospitalization rate associated with the 2nd dose of an mRNA vaccine (all ages, males & females) of 74.3% [42, Table 1]; we apply a case hospitalization rate of 80% for males 18–24, with 80% being approximately the average of 74.3% and the FDA rate of 86%
@@ 2.4. A Seventh Benefit Risk Scenario for mRNA-1273
The FDA’s BLA Clinical Review Memorandum [46] for mRNA-1273, completed 1/28/2022, describes a benefit-risk scenario that does not match any of the six scenarios appearing in the FDA’s published benefit-risk assessment of mRNA-1273 vaccination [9]. The BLA-Memo Scenario is described as follows in Section 4.7 of [46]:
The results of the modeling assessment predicted that in the population of adults at highest risk of vaccine associated myocarditis/pericarditis (males ages 18–25 years), vaccination of 1 million individuals with two primary series doses of mRNA-1273 vaccine would prevent 76,362 COVID-19 cases, 1,755 hospitalizations, 421 ICU visits and 4 deaths due to COVID-19, but would cause 148 myocarditis/pericarditis cases resulting in 128 hospitalizations, 47 ICU visits, and no deaths.
In Table 4 below, the BLA-Memo Scenario is compared to the FDA’s scenarios of [9, Table 2], in particular to Scenario 1, the “most likely scenario.” There are many noteworthy differences: e.g., dose-one VAM/P risk is included in the BLA-Memo Scenario but not in the scenarios of [9], and, in all scenarios of [9], the FDA predicts 0 ICU admissions owing to VAM/P while the BLA Memo predicts 1 million full vaccinations would cause 47.111111Subsequent analyses provide further evidence that the FDA underestimates VAM/P severity in [9]: A meta-analysis [47] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis with a mean length of stay (LOS) of 2.8 days and a 23.2% ICU-admission rate. Another pooled analysis [48] of 39 studies among 129 adults age 18 years or older (mean age 24) reported a median LOS of 3.0 days, 14% treated for heart failure, 8.1% admitted to the ICU, and 2 deaths (one after mRNA-1273 vaccination). A third meta-analysis [49] included 218 cases with a median LOS of 5.8 days with 3 deaths reported (one following mRNA-1273 vaccination), but provided no ICU-admission rate.
{adjustwidth}
-0.3cm
Table 4. FDA scenarios projecting benefit-risk outcomes per million 18–25-year-old males vaccinated with two primary series doses of Moderna RNA-1273.*
@@
* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2. Outcomes for the BLA-Memo scenario are from Section 4.7 of [46]. Outcomes for Scenarios 1–6 are from Table 2 of the FDA publication [9] with Scenario 1 being the “most likely scenario.”
Consistent with the BLA Memo Scenario, the FDA’s 1/31/22 press release [12] announcing the full approval of the mRNA-1273 vaccine indicates some mRNA-1273 VAM/P cases do require ICU care. Both the BLA Memo [46, Section 4.7] and the FDA’s benefit-risk assessment of mRNA-1273 [9, Section 2.3.2.2] attribute their contradictory ICU risk estimates to VSD data. The BLA Memo lacks a VSD-data reference, while the FDA author team of [9] cite a source of VSD data [50] whose collection cutoff point is 21 August 2021, over four months before the team completed their mRNA-1273 benefit-risk analysis [9].
The dated VSD data from [50] upon which the FDA [9] relies is scant, reporting only 21 postvaccination myo/pericarditis cases for those 18–29 (males and females, BNT162b2 and mRNA-1273, doses 1 and 2), with 18 of the cases ($\approx 86\%$) requiring hospitalization [50, p. 16]. For comparison, note that Buchan et al.  report a total of 74 myo/pericarditis cases after the 2nd dose of mRNA-1273 for the age range 18–29 based on Ontario data from 6/1/21 to 9/4/21 [42, Table 4]. Moreover, all these Ontario cases occurred during a 7-day risk-window postvaccination while the 21 cases reported in [50, p.  16] occurred in a 21-day risk-window postvaccination. A comparison of Public Health Ontario’s myo/pericarditis data 6/1/21– 9/4/21 [42, Table 4] to VSD data through 10/9/21 [10, pp. 8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million second doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system—see Appendix 3 of Supplement 2.
Because the BLA Memo’s quantification of the benefits and risks of mRNA-1273 vaccination does not closely relate to any of the scenarios of the FDA’s paper [9], it is unclear what role the analysis of [9] played in the FDA’s regulatory approval process for mRNA-1273. It is noteworthy that the FDA’s projection of no VAM/P-related ICU admissions for mRNA-1273 appears in a paper [9] submitted 5/19/22 well over three months after the appearance of FDA documents [12] and [46], with the former indicating ICU care is needed for some mRNA-1273 VAM/P cases and the latter projecting 47 VAM/P ICU admissions per million mRNA-1273 vaccinations of males 18–25.
The FDA’s published benefit-risk assessment [8] of Pfizer’s BNT162b2 omits ICU-admissions data for VAM/P. Funk et al. state in the last paragraph of the introduction of [8], ”Vaccine attributable myocarditis/pericarditis cases, hospitalizations, intensive care unit (ICU) admissions, and deaths were selected to pair with benefit endpoints for benefit-risk comparison.” However, in Section 2.3.1, Funk et al. indicate there are three risk endpoints, omitting ICU-admissions. There were, in fact, myo/pericarditis-related ICU admissions following BNT162b2 vaccination. The FDA acknowledges this in its 8/23/21 press release [11] announcing the full approval of BNT162b2 (Comirnaty). Moreover, VSD data collected through 8/21/21 and presented at the CDC’s ACIP meeting, August 2021, indicates that for the 12–17 age group there were 19 hospitalizations for post-BNT162b2 myo/pericarditis and 36.8% (7 of the 19) hospitalizations involved an ICU admission [50, p. 16].
2.5. Summary
In [9], the FDA substantially overestimated the hospitalizations-prevented benefit of mRNA-1273 vaccination for 18–25-year-old males by assuming COVID-hospitalization risk for this age group is equivalent to that for the 26–35 and 36–45 groups, by including incidental admissions, by ignoring prior-infection protection, and by assuming in its most likely scenario a COVID-19 case rate so high that the corresponding infection rate for the unvaccinated would be essentially 100% over 5 months. The FDA also substantially underestimated hospitalization risk for VAM/P, with available evidence indicating risk to be well over twice that of its most likely scenario—see Table 3 above. ICU-admission risk for VAM/P was underestimated as well (confirmed by FDA documents [46] and [12]).

### v3

@@ 2.2. The FDA’s mRNA-1273 Benefits Assessment
$\bullet$ The FDA assumed in its most likely scenario (as well as Scenarios 5, and 6) that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months. Throughout the pandemic, the number of COVID infections has always been significantly larger than the corresponding number of reported cases. For the period February 2020 through September 2021, the CDC estimated the number of infections to be four times the corresponding number of reported cases [25]. Moreover, the case-to-infection multiplier likely increased (e.g. owing to increased home testing) during the Omicron-dominant period in the U.S., which started mid-December 2021. If we conservatively assume that the CDC’s 4:1 infection:case ratio holds over the evaluation period, we find the FDA’s evaluation-period case-count projection of 274,947 per 1 million unvaccinated 18–25-year-old males (Scenarios 1, 5, 6) corresponds to 4 · 274,947 = 1,099,788 infections, well over 1 million. Rounding down, we deem the FDA’s case-total assumption to be equivalent to the assumption that among one million unvaccinated males, there would be, over 5 months, 1,000,000 infections. Having well over one million infections in a group of 1 million 18–25 year-olds over 5 months is implausible. Multiple infections over a five-month period would be unlikely given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [37]). [^4] The preprint [37], posted 6/10/21, was later published [38]. Even direct inoculation with SARS-CoV-2 does not necessarily produce infection: in the challenge study [39] of the 36 inoculated, 16 “remained uninfected.” Thus, for 3 of the FDA’s 5 Omicron-based scenarios, including its most likely scenario, the FDA essentially assumed that every unvaccinated male 18–25 would be infected by COVID over the evaluation period.
@@ 2.3. The FDA’s mRNA-1273 Risks Assessment
The FDA relied on its Biologics Effectiveness and Safety (BEST) System for the VAM/P case rates appearing in Table 1 above and used the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of VAM/P cases requiring hospitalization [11, Section 2.3.2].
VAM/P-case rates vary considerably depending on the case definition employed and the extent of stratification of risk, based on age, sex, vaccine manufacturer and dose, with most cases occurring postdose 2 [9]. For its mRNA-1273 assessment, the FDA four BEST-System data partners reported, with wide confidence intervals, the following myo/pericarditis incidence rates for 18–25-year-old males during the 7-day interval following dose 2: approximately 35, 70, 146, and 256 cases per million second doses [40, p, 12]. The FDA’s BEST System uses health-claims data not validated by a complete review of patients’ medical charts [11, Section 4], and has “inherent limitations, such as small sample sizes and imperfect sensitivity of ICD-10 codes to identify these rare outcomes” [10, Section 4]. We rely on four sources of data to estimate the VAM/P rate for dose 2 among 18–25-year-old males: (i) Buchan et al. [41] who apply the Brighton Collaboration’s case definition, (ii) Patone et al. [42] who consider only hospitalized cases of myocarditis, (iii) Sharff et al. [43] who apply the CDC’s case definition [44], and (iv) the FDA (with VSD-based adjustments).
The FDA underestimated mRNA-1273 VAM/P risk:
@@
$\bullet$ The FDA omitted dose-one VAM/P events. For its mRNA-1273 analysis [11], the FDA computed the evaluation-period benefits of two doses of mRNA-1273 without accounting for the VAM/P risk associated with dose 1, potentially biasing the FDA’s analysis in favor of vaccination. In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [10, Section 2.3.1]. Hence, for BNT162b2, the FDA did account for 1st-dose VAM/P risk but did not assess the benefits of dose 1 alone during the interval starting with dose 1 and ending at the point of full vaccination, potentially biasing the analysis against vaccination. Because adverse events of vaccination other than VAM/P were disregarded by the FDA and mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit-risk analysis assessing benefits only postdose 2, the choice to include 1st-dose VAM/P risk is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in the benefit-risk analysis of its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 2.4. Thus, we include 1st-dose VAM/P risk for mRNA-1273 in our main analysis. In Supplement 2, we conduct a sensitivity analysis showing that accounting for dose 1 benefits does not substantially change our benefit-risk conclusions.
@@
$\bullet$ The FDA disregarded evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2. In a “Review Memorandum Addendum” [45], dated 11/19/21, the FDA describes multiple sources of evidence that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2, including reports from regulatory agencies of five countries, data from Ontario later analyzed by Buchan et al. [41], and data from the CDC: “[T]he most recent data presented by the Center for Disease Control and Prevention (CDC) analyzed from the Vaccine Safety Datalink (VSD) indicates a higher rate of myocarditis with the Moderna COVID-19 Vaccine than the Pfizer-BioNTech COVID-19 Vaccine ….”
@@
The “most recent data” from the VSD to which the FDA refers was presented at the CDC’s October 2021 ACIP meeting [46]. Adjusted-rate ratios for 18–39-year-old males from the table on page 27 of [46] suggest the 2nd-dose VAM/P rate for mRNA-1273 is 2.26 times that of BNT162b2. Data from other sources suggests higher rate ratios e.g,, a 2nd-dose rate ratio of 5.1 was reported by Buchan et al. for males 18–24 [41, p. 5] and of 5.2 was reported by Husby et al. for males 12–39 [47, paragraph after Figure 3]. A study by Patone et al. of myocarditis cases requiring hospitalization within a 28-day risk interval after a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and 12 per million 1st doses, with corresponding BNT162b2 rates being 12 per million 2nd doses and 3 per million 1st doses [42, p. 4].
@@
Based on “preliminary findings” [45, p. 2], having “large uncertainty” owing to “small numbers of observed events” [40, p. 16], as well as omitting dose-1 cases, the FDA suggests that for 18–25-year-old males VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [11, Table 4] vs. 131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [10, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations. Throughout this paper, we assume that VAM/P rates for males 18–24 well approximate those for males 18–25.
Using the rate ratios and BNT162b2 VAM/P-risk assessment discussed above, we compute in Section 6.1 of Supplement 1 a VAM/P hospitalization rate of 224 per million 2nd doses of mRNA-1273 for males 18–25 years old.
The FDA’s BEST System also provided evidence that VAM/P risk for mRNA-1273 might be over twice that for BNT162b2. An FDA briefing document for a meeting held 10/14/2021 indicates that one of the agency’s BEST datapartners reported a postvaccination myo/pericarditis incidence rate of 283.7 (95% CI 145.2-573.5) cases per million full mRNA-1273 vaccinations (7 day risk window after each dose) [48, Section 2.3.3]. Applying a background rate of 4 cases per million per week for 18–25-year-old males [Supplement 1, Section 6.3] to each dose, we find the VAM/P case rate corresponding to 283.7 is 275.7 per million full vaccinations. Assuming a case-hospitalization rate of 100%, as in FDA’s BNT162b assessment, we obtain 275.7 VAM/P hospitalizations per million full mRNA-1273 vaccinations vs. 131 per million full BNT162b vaccinations.
$\bullet$ The FDA’s meta-analysis of its BEST-System data [11, Section 2.3.1] yielded VAM/P-rate estimates substantially lower than those of major international studies by Buchan et al. [41] and Patone et al. [42], as well as a U.S. study by Sharff et al. [43] [^5] The conclusion of [43] includes “We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm, which depends on select hospital discharge diagnosis codes. The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees.” . In Table 3 below, we compare the FDA’s projected mRNA-1273 VAM/P case and hospitalization rates to those of the other sources discussed above. (We have excluded VAM/P-rate sources, such as [49], depending on strictly passive surveillance—see Section 6.8 of Supplement 1.)
@@
Table 3. Estimated VAM/P Events Per Million 2nd Doses of mRNA-1273 Among Males 18–24 or 18–25 years old.a Based on Data Available Before 1 January 2022
@@
a For discussions of the data in columns 3–6, see Section 6 of Supplement 1; in particular, data in column 4 are estimates based on findings from [42] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [41]. To obtain the column-5 VAM/P case rate of 525 per million we applied the background myo/pericarditis rate suggested by [41].
b Buchan et al. applied the Brighton Collaboration’s case definition for myo/pericarditisx, which reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [41, p. 5].
c This estimated VAM/P-case rate is irrespective of the corresponding 1st-dose type. For homologous dosing (1st dose also mRNA-1273), Table 3 of [41] indicates the rate rises to at least 320 per million when the 2nd dose is taken less than 8 weeks after the 1st (with 4–6 weeks being the recommended interval [50]).
d For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [10, Section 2.3.2.2]. For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [11, Section 2.3.2.2]. We apply the 86% rate in columns 4–6 (which, for column 5, matches the case hospitalization rate: 6 of the 7 cases were hospitalized). For column 3, Buchan et al. report a case-hospitalization rate postdose 2 of an mRNA vaccine (all ages, males & females) of 79.2% ($\approx 164/207\times 100$%) [41, Table 1]; we expect the rate to be higher for the age range 18–24, but, conservatively, choose a hospitalization rate of 80%.
@@ 2.4. A Seventh Benefit Risk Scenario for mRNA-1273
The FDA’s BLA Clinical Review Memorandum [51] for mRNA-1273, completed 1/28/2022, describes a benefit-risk scenario not matching any scenario of the FDA’s published benefit-risk assessment of mRNA-1273 [11]. The projections of the BLA-Memo Scenario [51, Section 4.7] are described in the initial row of data of Table 4 below and compared to the FDA’s scenarios of [11, Table 2] in particular to Scenario 1, the most likely scenario. Differences include the BLA-Memo Scenario’s accounting for 1st-dose VAM/P risk and projecting 1 million full vaccinations would cause 47 ICU admissions (vs. 0 admissions). [^6] Subsequent analyses provide further evidence that the FDA may underestimate VAM/P severity in [11]; e.g., a meta-analysis [52] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis and a 23.2% ICU-admission rate.
Table 4. FDA scenarios projecting benefit-risk outcomes per million 18–25-year-old males vaccinated with two primary series doses of mRNA-1273.*
@@
* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2. Projections for the FDA’a Scenarios 1–6 are from [11, Table 2]. BLA-Memo-scenario projections are from [51, Section 4.7].”
Consistent with the BLA Memo Scenario, the FDA’s 1/31/22 press release [12] announcing the full approval of the mRNA-1273 vaccine indicates some mRNA-1273 VAM/P cases do require ICU care. Both the BLA Memo [51, Section 4.7] and the FDA’s mRNA-1273 benefit-risk assessment [11, Section 2.3.2.2] attribute their contradictory ICU risk estimates to VSD data. The BLA Memo lacks a specific reference, while the FDA in [11] cites a source of VSD data [53] with collection cutoff point 8/21/2021, over four months before completion of the FDA’s mRNA-1273 assessment.
The VSD data from [53] cited by the FDA [11] is scant, reporting only 21 postvaccination myo/pericarditis cases for those 18–29 [53, p. 16]. A comparison of Buchan et al.’s myo/pericarditis data from Ontario [41, Table 4, collected 6/1/21– 9/4/21] to VSD data through 10/9/21 [46, pp. 8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million second doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system [Supplement 2, Appendix 3].